OpenStudy (anonymous):

Find the domain of the composite function f & g f(x)=1/5-x g(x)=6/x

OpenStudy (australopithecus):

f(g(x)) and g(f(x))

OpenStudy (australopithecus):

those are the two composites of the functions

OpenStudy (australopithecus):

do you understand how to solve the problem now?

OpenStudy (anonymous):

No i don't.

OpenStudy (australopithecus):

f(g(x))=1/5-g(x) f(6/x) = 1/5 - 6/x f(g(x)) = 1/5 - 6/x

OpenStudy (australopithecus):

understand now?

OpenStudy (australopithecus):

g(x) is just a different way of writing y it use to be that (x,y) now you will see (x,f(x))

OpenStudy (australopithecus):

we use this notation because it helps prevent confusion and allows us to talk about a point easily

OpenStudy (anonymous):

I still don't understand how you solve it any further than that.

OpenStudy (australopithecus):

f(x) = 1/5-x f(1) = 1/5- 1 f(2) = 1/5 - 2 f(3) = 1/5 - 3 f(555) = 1/5 - 555 Therefore, knowing g(x) = 6/x f(g(x)) = 1/5 - 6/x

OpenStudy (australopithecus):

we are just subing one function into another I cant explain it any better than that

OpenStudy (australopithecus):

with math problems just look at what is changing and try to use logic to figure out why it is changing or at least recognize the pattern in the change so you can mimic said change

OpenStudy (anonymous):

how is that the Domain?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

Ya the question was to find the domian

OpenStudy (australopithecus):

oh its asking for the domain of the composite of the function sorry didnt read it

OpenStudy (anonymous):

lol no wonder i was confused

OpenStudy (australopithecus):

f(g(x)) = 1/5 - 6/x well REMEMBER You cannot divide by zero

OpenStudy (australopithecus):

so you should know what value of x is not present in the domain

OpenStudy (australopithecus):

do you know set notation?

OpenStudy (anonymous):

ummm no

OpenStudy (australopithecus):

it would be (-infinity, 0),(0, +infinity) notice how I used round brackets on the 0 this is to say that everything infinitely close but not 0 is included in the domain, if I used a square bracket it would mean zero is included in the set which it is not Note: you always use round brackets with infinity

OpenStudy (anonymous):

oh ok.

OpenStudy (anonymous):

so is that the answer?